Suppose you play a game with your friends. You tell your friends to each pick a number, and then you tell them to:
1.Multiply their number by 5.
2.Subtract 9 from that number.
3.Divide that number by 2.
Once they do that, they each tell you their number.
a. Ricky tells you his "new" number is -17. What number did Ricky start with?
b. Nina tells you her "new" number is 26.75. What number did Nina start with?
c. Isaac tells you his "new" number is 268. What number did Isaac start with?

Answer by Boreal(15235)

x is the number they started with
(5x-9)/2=-17
5x=-25
x=-5
5x-9=53.5, doubling 26.75
x=12.5
5x-9=536
5x=545
x=109